OpenStudy (anonymous):

For #1-4, state the amplitude, period, intervals, phase shift, vertical shift, domain, range and the first two positive asymptotes. y = tan(2x + pi) For #1-4, state the amplitude, period, intervals, phase shift, vertical shift, domain, range and the first two positive asymptotes. y = tan(2x + pi) @Mathematics

OpenStudy (anonymous):

rewrite tan=sin/cos will make it easier to solve

OpenStudy (anonymous):

huh? im not sure what you mean...can you write it out pls

OpenStudy (anonymous):

its a lot of work to solve this problem, ill give you some work efficient hints

OpenStudy (anonymous):

but can you atleast help me solve the range

OpenStudy (anonymous):

and intervals

OpenStudy (anonymous):

i have to switch computers right quick

OpenStudy (anonymous):

hey im back, are you there

OpenStudy (anonymous):

domain is \[(-\infty, \infty)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im trying to remember this all they way from when I took precalc sorry it will take me a while =( but im doing it right now

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok so let they that the formula is ->[a tan(bx-c)], to find domain its always (-infinity, infinity) range is (-a, a) period is (2pi/b) and finally to find phase shift its (c/b)

OpenStudy (anonymous):

wait i thought the phase shift was just c

OpenStudy (anonymous):

so the period would be (pi) and range (-1, 1) and phase (pi/2)

OpenStudy (anonymous):

nope its (c/b)

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

what about the range

OpenStudy (anonymous):

like i said its (-1,1)

OpenStudy (anonymous):

i mean how did you find it

OpenStudy (anonymous):

(-a, a)

OpenStudy (anonymous):

sin/cos=1/1=1

OpenStudy (anonymous):

ok thanks for the help